OpenStudy (anonymous):

Find the equation of the tangent plane to the surface determined by (x^4*y*2)+z−20=0 at x=2, y=2. z=

OpenStudy (anonymous):

i think this looks like example 1 here http://tutorial.math.lamar.edu/Classes/CalcIII/TangentPlanes.aspx

OpenStudy (anonymous):

we can try it if you like

OpenStudy (anonymous):

start with \[z=20=x^4y^2\] that is the equation right? then at x = 2, y = 2 you get \[z = 20-64=-44\] \[f_x=-4x^3y^2\] and \[f_y=-2x^4y\] at x = 2, y = 2 you get \[f_x=-128\] \[f_y=-64\] and the plane will be \[z+44=-128(x-2)-64(y-2)\] and then some algebra should complete it

OpenStudy (anonymous):

of course i could be way wrong, so i wouldn't bet money on this

OpenStudy (anonymous):

Thats exactly correct! Thanks for the help

OpenStudy (anonymous):

I appreciate the work you showed so I could follow the steps